supply function indicates the number of units of a commodity that suppliers are
willing to bring to the marketplace as a function of the price consumers are willing to pay. The following function is the supply function 9s = 0.5p2 – 200 where qs equal
no. of units supplied (stated in thousands) and p equals the selling price. a. What class of function is this? b. What quantity would be supplied if the market price is $30? $50? c. What price would result in 0 units being brought to the marketplace?
1
Expert's answer
2021-09-09T09:01:10-0400

(a) quadratic equation.

(b) "qs=0.5p^2-200"

At p= 30,

"qs= 0.5(30^2)-200=250"

At p=50,

"qs=0.5(50^2)-200=1050"

(c)

"qs=0.5p^2-200"

"0= 0.5p^2-200"

"p^2=400"

"p=20"